OpenStudy (loser66):

find partial solution of x'= Ax + b where \(A=\left[\begin{matrix}2&4\\-2&-2\end{matrix}\right]\) and \(b=\left[\begin{matrix}8sin2t\\8cos2t\end{matrix}\right]\) Please, help

OpenStudy (loser66):

I have characteristic equation for A is r^2 +4 =0 therefore, r = \(\pm2i\) so, eigenvector for r=2i is \(\left[\begin{matrix}1+i\\-1\end{matrix}\right]\) and eigenvector for r = -2i is \(\left[\begin{matrix}1-i\\-1\end{matrix}\right]\)

OpenStudy (loser66):

and .... stuck!! :)

OpenStudy (loser66):

My net is so bad here, if you don't see me, it doesn't mean that I don't care about the help. I 'll be back right after I can. ( In case I am kicked out of the net) Thanks in advance.

OpenStudy (loser66):

@TuringTest

OpenStudy (turingtest):

man I'm rusty on DE systems, but my answer to representing 1 + i would be\[\cos(0)+i\sin(\frac\pi2)\]if you can use two different angles. I'd have to review DE's to know if that's allowed, so you tell me if that helps

OpenStudy (turingtest):

you need it in terms of t, huh...

OpenStudy (loser66):

so, to plug them in fundamental matrix, what should I do?

OpenStudy (turingtest):

that's where I'd have to review DE's, I feel like I'm on shaky ground with this...

OpenStudy (loser66):

I am waiting for you. Take time to review and teach me then, please, please, please. final is upcoming, I must master those stuff

OpenStudy (turingtest):

I have a few things to study myself, but it's worth looking over if you haven't tried this resource I'd recommend it; it's what I'll be using: http://tutorial.math.lamar.edu/Classes/DE/ComplexEigenvalues.aspx

OpenStudy (loser66):

Thanks for the link. I will.

OpenStudy (turingtest):

I know your eigenvectors are right, but I keep getting the wrong answer for the first eigenvector"\[(2-2i)\eta_1+4\eta_2=0\implies\eta_2=\frac12(i-1)\implies\vec\eta=\binom 2{i-1}\]I know it's stupid to ask you to find my mistake when I'm the one supposed to be helping you, but I don't feel confident moving on until I can at least catch up to where you are :P

OpenStudy (turingtest):

could somebody just point out my mistake so I can move on? I know it's an algebra error or something dumb....

OpenStudy (loser66):

@TuringTest I am sorry when not being here, I cannot ignore the Asker when I can do something for them.

OpenStudy (turingtest):

Oh it's fine, I just figured the message system isn;t working very well... I'm reading up on my own stuff in the meantime, no rush

OpenStudy (schrodingers_cat):

I got the same thing turing \[\left(\begin{matrix}1 \\ (i-1)/2\end{matrix}\right)\]

OpenStudy (turingtest):

yeah, but wolfram disagrees :(

OpenStudy (schrodingers_cat):

In the end you are going to have to use one of the four main methods for solving differential equations either diagonalization, undetermined coefficients, Laplace transforms or variation of parameters.

OpenStudy (turingtest):

I was thinking variation of paramaters would be easiest as well, but I was trying a different approach

OpenStudy (schrodingers_cat):

Also, on math portal i got the answer above whereas on wolfram gave me his answer :P.I'm sure they are just multiples of each other by some factor :P

OpenStudy (schrodingers_cat):

I choose for x to be i and simplified and got the answer same as on math portal.

OpenStudy (turingtest):

well, whatever my mistake is, or isn't, the next step is to get one of the solutions like so\[\vec x_1(t)=e^{2i}\binom{-i-1}1=[\cos(2t)+i\sin(2t)]\binom{-i-1}1\]we multiply this out and factor out the i to get an answer of the form\[\vec x_1=\vec u(t)+i\vec v(t)\]

OpenStudy (schrodingers_cat):

Yes that is correct, taking the real and imaginary parts of x1 will you a fundamental set of solutions if the wronskian is not 0.

OpenStudy (turingtest):

is the wronskian zero? I'm not sure how to check with systems...

OpenStudy (turingtest):

hoping that the wronskian isn't zero... that gives\[\vec x_1(t)=\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}+i\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}\]and, again assuming the wronskian isn't zero, these guys are linearly independent, so the general complimentary solution is\[\vec x_1(t)=c_1\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}+c_2\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}\]if people could checkme on that I'd appreciate it, I don't trust my algebra :P

OpenStudy (loser66):

ohohoh,,, that's what my book says, I just don't get why from those eigenvector they can get it

OpenStudy (turingtest):

oh I see a mistake, that should be sine on the second row in the second solution

OpenStudy (turingtest):

\[\vec x_1(t)=\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}+i\binom{-\cos(2t)-\sin(2t)}{\sin(2t)}\]which implies\[\vec x(t)=c_1\binom{-\cos(2t)-\sin(2t)}{\cos(2t)}+c_2\binom{-\cos(2t)-\sin(2t)}{\sin(2t)}\]

OpenStudy (turingtest):

yeah, @Schrodingers_Cat and I both got the same other answer... But wolfram agrees with yur book, so I'd stick with that for now

OpenStudy (schrodingers_cat):

This is what I got however I got sin(2t) is positive on the top row of of the real part of the solution because i(-i) =1 right?

OpenStudy (turingtest):

I didn't multiply by i, I just factored it out

OpenStudy (turingtest):

oh I see

OpenStudy (turingtest):

\[\vec x_1(t)=[\cos(2t)+i\sin(2t)]\binom{-i-1}1\\=\binom{-i\cos(2t)+\sin(2t)-\cos(2t)-i\sin(2t)}{\cos(2t)+i\sin(2t)}\\=\binom{-\cos(2t)+\sin(2t)}{\cos(2t)}+i\binom{-\cos(2t)-\sin(2t)}{\sin(2t)}\]did I fix it?

OpenStudy (schrodingers_cat):

Yep that looks good :) Now to find the wronskian you have to do this \[\det \left[\begin{matrix}-\cos(2t)+\sin(2t) & -\cos(2t) -\sin(2t) \\ \cos(2t)& \sin(2t)\end{matrix}\right]\]

OpenStudy (turingtest):

doesn't look like what I'd expect it to for a system, but I'll take your word for it :)

OpenStudy (turingtest):

well that's 1, so I guess we're good to continue...

OpenStudy (schrodingers_cat):

Yeah the Wronskian looks like this W(u,v)(t) = det[u,v]

OpenStudy (schrodingers_cat):

Yep it is one.

OpenStudy (turingtest):

I thought wronskian was W=det[x1 x2] [x1' x2'] which wouldn't look like that, but I have a feeling that they wouldn't make this prob;em any harder than it already is by throwing us a cruveball now, we still have the non-homogenious part to do :P

OpenStudy (turingtest):

my book says we can always just call the second matrix linearly independent... whatever I trust you and I want to press on XD

OpenStudy (schrodingers_cat):

Yeah your just trying to find if the vectors u and v in the solution x1 are multiples of eachother.

OpenStudy (schrodingers_cat):

So, next would be the method to choose to solve the nonhomgenous system

OpenStudy (turingtest):

good enough for me\[\vec x_c(t)=c_1\binom{-\cos(2t)+\sin(2t)}{\cos(2t)}+c_2\binom{-\cos(2t)-\sin(2t)}{\sin(2t)}\]is our complimentary is where we got last I think

OpenStudy (turingtest):

yep

OpenStudy (turingtest):

gotta read again lol

OpenStudy (turingtest):

I'm looking at this btw: http://tutorial.math.lamar.edu/Classes/DE/NonhomogeneousSystems.aspx

OpenStudy (schrodingers_cat):

I was thinking probably variation of parameters but the only problem with that is we have to find the inverse of the our fundamental which might isn't to frustrating hopefully :)

OpenStudy (schrodingers_cat):

We already know the determinant which is nice :)

OpenStudy (turingtest):

I agree, but either way we needed the complimentary, so I figured I'd get that out of the way before even thinking about what comes next. I'm taking a quick break, back in 5. Feel free to forge ahead :)

OpenStudy (loser66):

feel free to discuss and please, don't delete anything. I 'll be back to get. Thanks for discussing

OpenStudy (turingtest):

yeah variation of parameters is totally the way to go we need to get the particular formula with\[\vec x_p(t)=X\int X^{-1}\vec g(t)dt\]where \(X\) is the combination of the two solution vectors \([\vec x_1|\vec x_2]\) we can get the inverse \(X^{-1}\) pretty easily with the formula\[X=\left[\begin{matrix}a&b\\c&d\end{matrix}\right]\implies X^{-1}=\frac1{\det X}\left[\begin{matrix}d&-b\\-c&a\end{matrix}\right]\]

OpenStudy (turingtest):

we already found that \(\det X=1\) above, so I am hoping you can actually take it from here. Let me if you get stuck!

OpenStudy (schrodingers_cat):

Yep that is correct its like this problem was set up so the determinant was 1 :)

OpenStudy (schrodingers_cat):

lol

OpenStudy (turingtest):

at least they threw us one bone :P long problems, these types are....

OpenStudy (schrodingers_cat):

Yep also Loser66 its important that you learn the other methods as well because if have simple solutions to your characteristic equation diagonaliztion will save you alot of time especially on a test.

OpenStudy (loser66):

I got everything but the step from eigenvector form\(\left[\begin{matrix}-1-i\\1\end{matrix}\right]\) to cos, sin, sin cos,.... If I past that step. I am ok with the rest.

OpenStudy (turingtest):

the solution to your characteristic equation is \(e^{rit}\), use Euler's formula...

OpenStudy (loser66):

I know how to take wronskian, apply Cramer's rule to get u' t and then take integral to get u (t) , then plug back to get partial, but the so called Euler's formula

OpenStudy (turingtest):

or rather as an eigenvalue \(\large e^{\lambda it}\) Euler's formula, which you absolutely *must* know in any DE class, is\[\large e^{i\theta}=\cos\theta+i\sin\theta\]

OpenStudy (loser66):

we have \(\lambda = \pm2i \) that mean e^0 (cos 2t + sin 2t) , then?? do what to eigenvectors to get fundamental matrix?

OpenStudy (schrodingers_cat):

Yes remember loser66 that a general solution to system of differential equations takes the form \[x = c_{n}e^{\lambda t}\] The problem is evaluating the exponential if you have imaginary solutions Using Euler's formula as turing stated above allows us to do this!!!!!!!!!!!!!!!!!!!!

OpenStudy (turingtest):

it can be shown that if the roots to the characteristic or eigenvalue is complex, \[\lambda=\mu\pm\theta i\]we can write the solution set as\[\large y_{1,2}(t)=e^{(\mu\pm\theta i)t}=e^{\mu t}(\cos(\theta t)+i\sin(\theta t) )\]from which it can be shown that this can be written\[\large y(t)=c_1e^{\mu}\cos(\theta t)+c_2\sin(\theta t)\]

OpenStudy (schrodingers_cat):

Oops there should be an eige vector in there after the Cn :)

OpenStudy (turingtest):

Yeah, my bad thanks. knew I was bound for some typos it can be shown that if the roots to the characteristic or eigenvalue is complex, \[\lambda=\mu\pm\theta i\]we can write the solution set as\[\large y_{1,2}(t)=e^{(\mu\pm\theta i)t}=e^{\mu t}(\cos(\theta t)+i\sin(\theta t) )\]from which it can be shown that this can be written\[\large y(t)=c_1e^{\mu t}\cos(\theta t)+c_2e^{\mu t}\sin(\theta t)\]

OpenStudy (schrodingers_cat):

Oh no I was talking about the one I posted above lol

OpenStudy (turingtest):

oh mine has typos too :P speaking generally above here, not just about systems^

OpenStudy (turingtest):

in any event, Loser66 the point is we just got one eigenvector and multiplied it by the complimentary solution to the characteristic... erm is that the right term we multiplied one of the eigenvectors by \[\large e^{\lambda t}=e^{(\mu\pm\theta) t}=\cos(\theta t)+i\sin(\theta t)\]

OpenStudy (loser66):

oh, I got it. thanks a lot

OpenStudy (turingtest):

cool, welcome!

OpenStudy (turingtest):

and thanks @Schrodingers_Cat too!

OpenStudy (schrodingers_cat):

Yep, hope this helps good luck :)

OpenStudy (loser66):

Thanks for your time.